# Efficient Multi-scale Network with Learnable Discrete Wavelet Transform for Blind Motion Deblurring

### 29 Dec 2023 v1

@@ Abstract
Coarse-to-fine schemes are widely used in traditional single-image motion deblur; however, in the context of deep learning, existing multi-scale algorithms not only require the use of complex modules for feature fusion of low-scale RGB images and deep semantics, but also manually generate low-resolution pairs of images that do not have sufficient confidence. In this work, we propose a multi-scale network based on single-input and multiple-outputs(SIMO) for motion deblurring. This simplifies the complexity of algorithms based on a coarse-to-fine scheme. To alleviate restoration defects impacting detail information brought about by using a multi-scale architecture, we combine the characteristics of real-world blurring trajectories with a learnable wavelet transform module to focus on the directional continuity and frequency features of the step-by-step transitions between blurred images to sharp images. In conclusion, we propose a multi-scale network with a learnable discrete wavelet transform (MLWNet), which exhibits state-of-the-art performance on multiple real-world deblurred datasets, in terms of both subjective and objective quality as well as computational efficiency. Our code will be open-sourced on github later.
${}^{\ast}$ Equal contribution
${}^{{\dagger}}$ Coresponding Author
@@ 1 Introduction
Figure 1: Performance comparison on the RealBlur-J [23] test dataset in terms of PSNR and GMACs. Our proposed MLWNet achieves both higher accuracy and smaller complexity in comparison with other state-of-the-arts.
@@
The coarse-to-fine algorithm has another inherent defect. During progressive restoration, the solution of the upper-level problem takes the solution of the lower-level as initialization [9]. Although this reduces the difficulty of solving the upper-level problem, due to the smaller resolution of lower-level spatial, the features transmitted upwards are semantically precise but spatially ambiguous, thus the reatoration ability of multi-scale network in spatial details is limited. Hence, there is an urgent need to improve the quality of restoration of high-frequency detail part. A simple approach is to introduce a frequency domain transform as an alternative to a spatial domain transform. This would provide the algorithm with a choice and direct its attention to different frequencies. In this paper, we consider using the discrete wavelet transform for the following reasons:
@@ 2 Related Work
Multi-scale Deblurring Algorithm. Multi-scale motion blur removal methods aim to achieve progressive clear image recovery using a multi-stage or multi-sub-network approach. The multi-scale algorithm previously used in the field of image deblurring is based on MIMO [20, 27, 21, 3, 13]. Recently, Cho et al. [3] proposed MIMO-UNet, which employs a multi-output single decoder as well as a multi-output single decoder to simulate a multi-scale architecture consisting of stacked networks, thereby greatly simplifying complexity. Zamir et al. [37] designed a multi-stage progressive image recovery network to simplify the information flow between multiple sub-networks. As shown in Sec. 3.1, our proposed multi-scale approach further simplifies the structure of multi-scale networks and better balances the accuracy and complexity.
Application of Frequency in Deblurring. Frequency is an important property of images, in recent years frequency domain has been introduced into numerous DNNs to be solved for various tasks [25, 32, 41, 10, 35, 22]. Mao et al. [19] proposed DeepRFT, which introduced a residual module based on the Fourier transform into an advanced algorithm. Kong et al. [10] proposed FFTformer, which introduced the Fourier transform into a feed-forward network to effectively utilize different frequency information. Zou et al. [42] proposed SDWNet, which utilized frequency domain information as a complement to spatial domain information. Different from these methods, we introduce a learnable 2D-DWT capable of adapting to the data distribution and feature space, more suitable not only for dealing with digital images rich in mutation signals, but also for dealing with real blurring.
@@ 3.1 Multi-scale Baseline
For ease of understanding, we will introduce here the Encoder phase, multi-scale semantic Fusion phase (hereafter referred to as Fusion phase), and Decoder phase sequentially. As shown in Fig. 2, the Encoder stage consists of several gray Simple Encoder Blocks(SEB), where information flows top-down for feature extraction; the Fusion stage consists of several green Wavelet Fusion Blocks(WFB), where information flows bottom-up for semantic fusion at different scales; and the Decoder stage consists of several purple Wavelet Head Blocks(WHB), and the bottom-up flow of information enables gradual restoration.
@@
$E_{out}^{i}$ =$=\left\{\begin{array}[]{lr}\phi_{i}\left(embed(x)\right),&\text{if }i=i_{max}\\ \phi_{i}\left(E_{out}^{i-1}\right),&otherwise\\ \end{array}\right.$ (1)
where $\phi_{i}$ represent the block of corresponding layer of Encoder. The Fusion stage adapts and fuses information from different scales of semantics produced in Encoder to generate intermediate representations with deep semantics and shallow details. The Fusion stage’s outputs $E_{out}^{i}$ to the block of $i-th$ layer is expressed as in Eq. 2, where $\delta_{i}$ represents the block of the corresponding layer of Fusion.
$F_{out}^{i}$ =$=\left\{\begin{array}[]{lr}\delta_{i}\left(E_{out}^{i}+E_{out}^{i-1}\right),&\text{if }i=i_{min}\\ \delta_{i}\left(F_{out}^{i-1}+E_{out}^{i}\right),&otherwise\\ \end{array}\right.$ (2)
The Decoder stage utilizes the information passed by both the Encoder and Fusion stages to progressively upsample the feature maps and generating pre-output feature maps at each scale separately. Output $D_{out}^{i}$ to the block of the $i-th$ layer is denoted as in Eq. 3. Here $\xi_{i}$ represents the block of the corresponding layer of the Decoder.
$D_{out}^{i}$ =$=\left\{\begin{array}[]{lc}\xi_{i}\left(E_{out}^{0}\right),&\text{if }i=i_{min}\\ \xi_{i}\left(D_{out}^{i-1}\right),&\text{if }i>i_{min}\text{ and }i<i_{max}\\ \xi_{i}\left(D_{out}^{i-1}+F_{out}^{i-1}\right),&\text{if }i=i_{max}\\ \end{array}\right.$ (3)
@@ 4.1 Datasets and Implementation Details
We evaluate our method on the realistic datasets RealBlur [23] and RSBlur [24], as well as the synthetic dataset GoPro [20], they are all composed of blurry-sharp image pairs. During training we use AdamW optimizer ($\alpha=0.9$ and $\beta=0.9$) for a total of 600K iterations. The initial value of the learning rate is $10^{-3}$ and is updated with cosine annealing schedule. The patch size is set to be 256 × 256 pixels and we followed the training strategy of [38, 28]. For data augmentation, we only use random flips and random rotations. Our experiments are performed on NVIDIA GeForce RTX 3090 GPUs and Intel Xeon Gold 6130 CPU.
@@ 4.2 Comparison with State-of-the-Art Methods
Evaluations on the RealBlur dataset. Since our proposed method is driven by real-world deblurring, we first conduct comparisons on RealBlur [23]. The quantitative analysis results are shown in Tab. 1, the PSNR value of our method is 1.02dB and 0.49dB higher than state-of-the-art GRL on the RealBlur-J and RealBlur-R datasets respectively. Compared to other multi-scale architectures, our method has fewer model computational and running time while the performance is better. Fig. 4 shows visual comparison on RealBlur-J dataset, our method obtains restored text with richer edge details, while achieving the best contrast, sharpness, brightness, and structural details of the image.
@@ 5 Analysis and Discussion
Single-scale vs. Multi-scale. Given that each node of the baseline network is composed of SEB, we have deformed the multi-scale and single-scale according to the architecture. Multi-scale architecture has improved PSNR and SSIM values to varying degrees, indicating that the SIMO strategy we adopt helps remove motion blur to a certain extent. As shown in Sec. 3.1, we use SIMO for coarse-to-fine restoration, and only use multiple outputs to calculate multi-scale losses during training.
WFB and WHB. We default to using WFB and WHB via $\mathcal{L}_{wavelet}$ statute in this section, we also proved the effectiveness of LWN, because LWN is a subpart of the first two. After applying them, the PSNR index improved by 0.25dB compared to the multi-scale baseline, which shows that LWN and its extension modules can easily help multi-scale algorithms improve detail recovery capabilities. Fig. 7 shows the four types of feature maps generated after forward learnable wavelet convolution. We can see that the high and low frequencies of the input feature map are mixed together, and after wavelet group convolution, the network exerts different attention on the feature information in different directions or frequencies.
@@
SIMO WFB WFB $\mathcal{L}_{wavelet}$ PSNR SSIM MACs(G)
32.29 0.924 19.24
@@
Limitations and analysis. It is observed from Tab. 1-3 that our method fails to achieve the same expected high performance with synthetic blur as it does with realistic blur. We try to analyze the reasons for this phenomenon via dataset comparison and experimental results. First, the average frame synthesis method used in the synthetic blur dataset leads to unnatural discontinuous blur trajectories, thereby introducing strange high-frequency information interference (see Fig. 8). Second, synthetic data tends to produce an unnatural mixture of frequencies. It is easy to mix with the average color area at the edge of the texture, causing the confusion of high-low frequency information.

### 13 Mar 2024 v2

@@ Abstract
Coarse-to-fine schemes are widely used in traditional single-image motion deblur; however, in the context of deep learning, existing multi-scale algorithms not only require the use of complex modules for feature fusion of low-scale RGB images and deep semantics, but also manually generate low-resolution pairs of images that do not have sufficient confidence. In this work, we propose a multi-scale network based on single-input and multiple-outputs(SIMO) for motion deblurring. This simplifies the complexity of algorithms based on a coarse-to-fine scheme. To alleviate restoration defects impacting detail information brought about by using a multi-scale architecture, we combine the characteristics of real-world blurring trajectories with a learnable wavelet transform module to focus on the directional continuity and frequency features of the step-by-step transitions between blurred images to sharp images. In conclusion, we propose a multi-scale network with a learnable discrete wavelet transform (MLWNet), which exhibits state-of-the-art performance on multiple real-world deblurred datasets, in terms of both subjective and objective quality as well as computational efficiency. Our code is available on https://github.com/thqiu0419/MLWNet.
@@ 1 Introduction
Figure 1: Performance comparison on the RealBlur-J [23] test dataset in terms of PSNR and GMACs. Our proposed MLWNet achieves superiority in comparison with other state-of-the-arts.
@@
The coarse-to-fine algorithm has another inherent defect. During progressive restoration, the solution of the upper-level problem takes the solution of the lower-level as initialization [9]. Although this reduces the difficulty of solving the upper-level problem, due to the smaller resolution of lower-level spatial, the features transmitted upwards are semantically precise but spatially ambiguous, thus the restoration ability of multi-scale network in spatial details is limited. Hence, there is an urgent need to improve the quality of restoration of high-frequency detail part. A simple approach is to introduce a frequency domain transform as an alternative to a spatial domain transform. This would provide the algorithm with a choice and direct its attention to different frequencies. In this paper, we consider using the discrete wavelet transform for the following reasons:
@@ 2 Related Work
Multi-scale Deblurring Algorithm. Multi-scale motion deblur methods aim to achieve progressive recovery using a multi-stage or multi-sub-network approach. The multi-scale algorithm previously used in the field of image deblurring is based on MIMO [20, 27, 21, 3, 13]. Recently, Cho et al. [3] proposed MIMO-UNet, which employs a multi-output single decoder as well as a multi-input single encoder to simulate a multi-scale architecture consisting of stacked networks, thereby greatly simplifying complexity. Zamir et al. [37] designed a multi-stage progressive image recovery network to simplify the information flow between multiple sub-networks. As shown in Sec. 3.1, our proposed multi-scale approach further simplifies the structure of multi-scale networks and better balances the accuracy and complexity.
Application of Frequency in Deblurring. Frequency is an important property of images, in recent years frequency domain has been introduced into numerous DNNs to be solved for various tasks [25, 32, 41, 10, 35, 22]. Mao et al. [19] proposed DeepRFT, which introduced a residual module based on the Fourier transform into an advanced algorithm. Kong et al. [10] proposed FFTformer, which introduced the Fourier transform into a feed-forward network to effectively utilize different frequency information. Zou et al. [42] proposed SDWNet, which utilized frequency domain information as a complement to spatial domain information. Different from these methods, we introduce a learnable 2D-DWT capable of adapting to data distribution and feature space, more suitable not only for dealing with digital images rich in mutation signals, but also for dealing with real blurring.
@@ 3.1 Multi-scale Baseline
For ease of understanding, we will introduce here the Encoder phase, multi-scale semantic Fusion phase (hereafter referred to as Fusion phase), and Decoder phase sequentially. As shown in Fig. 2, the Encoder stage consists of several gray Simple Encoder Blocks(SEB), where information flows top-down for feature extraction; the Fusion stage consists of several green Wavelet Fusion Blocks(WFB), where information flows bottom-up for semantic fusion at different scales; and the Decoder stage consists of several purple Wavelet Head Blocks(WHB), and the bottom-up flow of information enables gradual restoration. Compared with WHB, SEB replaces LWN with a $1\times 1$ conv for channel scaling, and a $3\times 3$ depth-conv for feature extraction.
@@
$E_{out}^{i}$ =$=\left\{\begin{array}[]{lr}\phi_{i}\left(embed(x)\right),&i=i_{max}\\ \phi_{i}\left(E_{out}^{i-1}\right),&otherwise\\ \end{array}\right.$ (1)
where $\phi_{i}$ represent the block of corresponding layer of Encoder. The Fusion stage adapts and fuses information from different scales of semantics produced in Encoder to generate intermediate representations with deep semantics and shallow details. The Fusion stage’s outputs $F_{out}^{i}$ to the block of $i-th$ layer is expressed as in Eq. 2, where $\delta_{i}$ represents the block of the corresponding layer of Fusion.
$F_{out}^{i}$ =$=\left\{\begin{array}[]{lr}\delta_{i}\left(E_{out}^{i}+E_{out}^{i-1}\right),&i=i_{min}\\ \delta_{i}\left(F_{out}^{i-1}+E_{out}^{i}\right),&otherwise\\ \end{array}\right.$ (2)
The Decoder stage utilizes the information passed by both the Encoder and Fusion stages to progressively upsample the feature maps and generating pre-output feature maps at each scale separately. Output $D_{out}^{i}$ to the block of the $i-th$ layer is denoted as in Eq. 3.1. Here $\xi_{i}$ represents the block of the corresponding layer of the Decoder.
$D_{out}^{i}=\begin{cases}\xi_{i}\left(E_{out}^{i}\right),\qquad\qquad\qquad\qquad\quad i=i_{min}\\ \xi_{i}\left(D_{out}^{i-1}+F_{out}^{i}\right),\qquad i>i_{min}\text{ and }i<i_{max}\\ \xi_{i}\left(D_{out}^{i-1}+F_{out}^{i-1}+E_{out}^{i}\right),\qquad i=i_{max}\end{cases}$
@@ 4.1 Datasets and Implementation Details
We evaluate our method on the realistic datasets RealBlur [23] and RSBlur [24], as well as the synthetic dataset GoPro [20], they are all composed of blurry-sharp image pairs. During training we use AdamW optimizer ($\alpha=0.9$ and $\beta=0.9$) for a total of 600K iterations. The initial value of the learning rate is $10^{-3}$ and is updated with cosine annealing schedule. The patch size is set to 256 × 256 pixels and we followed the training strategy of [38, 28]. For data augmentation, we only use random flips and rotations. More experimental results can be found in the Appendix.
@@ 4.2 Comparison with State-of-the-Art Methods
Evaluations on the RealBlur dataset. Since our proposed method is driven by real-world deblurring, we first conduct comparisons on RealBlur [23]. The quantitative analysis results are shown in Tab. 1, the PSNR value of our method is 0.91dB and 0.49dB higher than state-of-the-art GRL on the RealBlur-J and RealBlur-R datasets respectively. Compared to other multi-scale architectures, our method has fewer model computational and running time while the performance is better. Fig. 4 shows visual comparison on RealBlur-J dataset, our method obtains restored text with richer edge details, while achieving the best contrast, sharpness, brightness, and structural details of the image.
@@ 5 Analysis and Discussion
Single-scale vs. Multi-scale. Given that each node of the baseline network is composed of SEB, we have deformed the multi-scale and single-scale according to the architecture. Tab. 5 shows that the multi-scale architecture improves the PSNR and SSIM values to varying degrees, proving that the SIMO strategy we employ helps to eliminate motion blur to some extent. As shown in Sec. 3.1, we use SIMO for coarse-to-fine restoration, and only use multiple outputs to calculate multi-scale losses during training.
WFB and WHB. We default to using WFB and WHB via $\mathcal{L}_{wavelet}$ statute in this section, we also proved the effectiveness of LWN, because LWN is a subpart of the first two. After applying them, the PSNR index improved by 0.25dB compared to the multi-scale baseline, which shows that LWN and its extension modules can easily help multi-scale algorithms improve detail recovery capabilities. Fig. 7 shows the four types of feature maps generated after forward learnable wavelet convolution. We can see that the high and low frequency information of the input feature map are mixed together, and after wavelet group convolution, the network exerts different attention on the feature information in different directions or frequencies.
@@
Method SISO MIMO SIMO
PSNR/SSIM 32.29/0.924 32.19/0.928 32.37/0.929
MACs(G) 19.24 21.83 19.29
Table 5: Comparison in various input and output modes.
SIMO WFB WHB $\mathcal{L}_{wavelet}$ PSNR SSIM MACs(G)
@@
Limitations and analysis. It is observed from Tab. 1-3 that our method fails to achieve the same expected high performance with synthetic blur as it does with realistic blur. We try to analyze the reasons for this phenomenon via dataset comparison and experimental results. First, the average frame synthesis method used in the synthetic blur dataset leads to unnatural discontinuous blur trajectories, thereby introducing strange high-frequency information interference (see Fig. 8). Second, synthetic data tends to produce an unnatural mixture of high and low frequencies. It is easy to mix with the average color area at the edge of the texture, causing the confusion of high-low frequency information.